OpenStudy (wade123):

i need someone to check my work! The position of an object at time t is given by s(t) = 1 - 10t. Find the instantaneous velocity at t = 10 by finding the derivative.

OpenStudy (wade123):

my answer: ds/dt=v s(t) = 1 - 10t ds/dt= d(1-10t)/dt d(1-10t)/dt = d(1)/dt - d(10t)/dt =0-10=-10

OpenStudy (wade123):

@sweetsunray

OpenStudy (johnweldon1993):

Indeed you are correct, since the derivative just comes out to be a constant (-10) that will be your instantaneous velocity at any time

OpenStudy (anonymous):

perfect wade!

OpenStudy (triciaal):

great job